OpenStudy (anonymous):

for F(x)=1/4x-5 find (fof^-1) (-7)

OpenStudy (anonymous):

 confirm this is ur problem please

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

i think the answer is 7/509 in order to find f'(x) you must change out x for y and solve for y

OpenStudy (anonymous):

scratch that - answer is -1/137

OpenStudy (anonymous):

F(x) = 1/(4x-5) so F-1(x) = 4x-5 i tested with x = 1 and they both came out to the same answer then in f(x) WHEREever you see x plug in 4x-5. plug in -7 and solvee

OpenStudy (anonymous):

ok, annd thank youu (: